OpenStudy (anonymous):

What is the solution of the system? Use elimination. 3x – 4y = –24 x + y = –1 A. (–4, 3) B. (0, 6) C. (3, 4) D. (4, 3) A, B, C, or D? Please explain (:

OpenStudy (anonymous):

3x – 4y = –24 EQ 1) x + y = –1 EQ2) mult by 4 then add

OpenStudy (anonymous):

3x – 4y = –24 4x + 4y = –4 add to eliminate y --------------- 7x =-28 x=-28/7 x=-4

OpenStudy (anonymous):

now sub x=-4 to EQ2) -4+y=-1 y=4-1 y=3 (-4,3)=(x,y) ...ans

OpenStudy (anonymous):

Thank you mark_o..

OpenStudy (anonymous):

yw...have fun solving..lol